OpenStudy (anonymous):

if n+2 is divisible by 2, is n: an even number? an odd number? or a prime number?

OpenStudy (kropot72):

N must be an even number to meet the requirement.

myininaya (myininaya):

\[\frac{n+2}{2}=integer\] let k be an integer Well test what happens if n=2k+1 this would imply n is odd \[\frac{2k+1+2}{2}\] We couldn't cancel the 2 on bottom :( ok so what happens if n=2k this would imply n is even so we would have \[\frac{2k+2}{2}=\frac{2(k+1)}{2}=k+1\] and k+1 is an integer since k is an integer

myininaya (myininaya):

The expression is divisible for one prime that prime being 2 since \[\frac{2+2}{2}=\frac{4}{2}=2\] But this prime is also even while all other primes are odd (and i'm not saying all odd are prime-lol; i love math)

myininaya (myininaya):

@artist do you have any questions?

OpenStudy (zepp):

What if our integer that we are looking for is \[\frac{n+2}{2}\] and that integer could be simplified as \[\frac{n}{2} + \frac{2}{2}\] Our lovely integer would be \[\frac{n}{2} + 1\] Since our answer is an integer, n must be an even number so it could be divisible by 2.

OpenStudy (asnaseer):

if (n+2) is divisible by 2, then (n+2) must be even. if you take 2 away from ANY even number, you are still left with an even number. therefore, since (n+2) is even, (n+2-2) must also be even. therefore n is even.

OpenStudy (anonymous):

If n is odd so is n+2. Primes are odd except 2 So n is always even